ROI, Margin, Turnover

Jarriot, Inc., presented two years of data for its Furniture Division and its Houseware Division.

Furniture Division:

Year 1 Year 2
Sales $33,870,000 $36,210,000
Operating income 1,422,540 1,520,820
Average operating assets 10,000,000 10,000,000

Houseware Division:

Year 1 Year 2
Sales $12,590,000 $13,139,800
Operating income 604,320 499,312
Average operating assets 5,000,000 5,000,000

Required:

Round the ROI and margin percentages to two decimal places (for example, enter the decimal .10555 as "10.56" percent). Round the turnover ratio to two decimal places.

1. Compute the ROI and the margin and turnover ratios for each year for the Furniture Division.

Furniture Division
ROI Margin Turnover
Year 1 % %
Year 2 % %

2. Compute the ROI and the margin and turnover ratios for each year for the Houseware Division.

Houseware Division
ROI Margin Turnover
Year 1 % %
Year 2 % %

3. Indicate the reason for change in ROI from Year 1 to Year 2 for the Furniture Division.

Indicate the reason for change in ROI from Year 1 to Year 2 for the Houseware Division.
Answer #1
1) Furniture division
ROI Profit Margin Turnover
Return On Assets = Net Income / Average Assets Profit Margin = Net Income / Sales Assets Turnover Ratio = Sales / average Assets
year1 = $1422540/10000000 = $1422540/33870000 = $33870000/10000000
=14.23 % =4.2 % =3.39 times
Year 2 = $1520820/10000000 = $1520820/36210000 = $36210000/10000000
=15.21 % =4.2 % =3.62 times
2) Houseware Division:
ROI Profit Margin Turnover
Return On Assets = Net Income / Average Assets Profit Margin = Net Income / Sales Assets Turnover Ratio = Sales / average Assets
year1 = $604320/5000000 = $604320/12590000 = $12590000/5000000
=12.09 % =4.8 % =2.52 times
Year 2 = $499312/5000000 = $499312/13139800 = $13139800/5000000
=9.99 % =3.8 % =2.63 times
3) Indicate the reason for change in ROI from Year 1 to Year 2 for the Furniture Division
Dure to increase of asset turnover.
Indicate the reason for change in ROI from Year 1 to Year 2 for the Houseware Division.
dure to decrease in profit margin ratio.